A turbine is spinning freely at 3750 rpm. It takes 10 min to coast to a stop due to friction in the bearings. How many revolutions doe the turbine make while stopping? Express your answer in revolutions.
Answer #1

Initial angular velocity wi = 3750rpm and final angular velocity wf = 0

Time t = 10mints

Angular distance = wi × t = 3750×10 = 37500 revolution.